OpenStudy (anonymous):

the derivative of an equation is 4x(x^2-4) it wants me to find the location of relative maximum, minimum and inflection points how do I do this?

OpenStudy (mrnood):

those points all are found when the derivative =0 so you need the followowing steps to start first factorise the term in brackets so you have (4x)(x-a)(x-b)

OpenStudy (freckles):

so we have \[y'=4x(x^2-4) \text{ for some } y=f(x)\] We can find the critical points f by setting y'=0

OpenStudy (anonymous):

they are 0 and 2 right

OpenStudy (freckles):

\[\text{ and } x=-2\]

OpenStudy (mrnood):

no - there are 3 points - just factorise as I said first

OpenStudy (mrnood):

or just let freckles post the answer for you

OpenStudy (anonymous):

2,0, and -2 okay then I do a sin chart?

OpenStudy (freckles):

well he had part of the answer

OpenStudy (freckles):

I think since it is a cal question he knows how to solve x^2-4=0

OpenStudy (freckles):

he just forgot or i could be wrong

OpenStudy (mrnood):

the way I would approach it is then to take the second derivative - i.e. differentiat the first derivative ....

OpenStudy (mrnood):

d (uv) = vdu + udv dx dx dx

OpenStudy (mrnood):

If the second derivative is -ve at the critical point then the point is a maximum If the second derivative is +ve at the critical point then the point is a minimum If the second derivative is 0 at the critical point then the point is a POI

OpenStudy (anonymous):

so I get the answer by taking the second derivative?

OpenStudy (mrnood):

take the second derivative f'' For each of the three x values you got above (0,2,-2) put them into the formula for the f'' and evaluate the value of f'' Then apply the rules in my last post to decide whether the point is a max, min or POI